OpenStudy (anonymous):

Quincy had a coupon for his favorite clothing store. He received a discount of eleven dollars on a pair of jeans that originally cost fifty-five dollars. Set up a percent proportion to determine the percent discount he received with his coupon. Which of the following is a denominator in the proportion? A. forty-four B. fifty-five C .thirty-five D. eleven HELP FREE MEDAL !

OpenStudy (jdoe0001):

I take the free medal, I mean is free, so free is free, as in free beer, right? :P

OpenStudy (anonymous):

Ahh man c'mon lol

OpenStudy (jdoe0001):

wait a second! so it's really not free =P

OpenStudy (anonymous):

for the questions answer its free

OpenStudy (jdoe0001):

$55 is the 100% of the original cost $11 is some % we dunno, say x% so $11 corresponds to $55 as 100% corresponds to x% \(\bf \cfrac{11}{x}=\cfrac{55}{100}\) solve for "x" :P

OpenStudy (anonymous):

gives me 20 is that the answer?

OpenStudy (jdoe0001):

x = 20% well, 20% = 20/100 = 1/5 is 11 1/5 of 55?

OpenStudy (anonymous):

so its 55?

OpenStudy (jdoe0001):

55 is the original amount you said you got 20, that is x = 20% that is 20% is 1/5 of the original amount so check yourself, is 11 one-fifth of 55?

OpenStudy (anonymous):

I guess?

OpenStudy (jdoe0001):

well, that's the idea, you're not supposed to guess you're suppose to calculate it

OpenStudy (anonymous):

yeah it is

OpenStudy (jdoe0001):

then I guess x = 20, or 20% so 11 is really 20% of 55 :)